A projectile motion us given by y=√3/3x - 4/3x^2 where (x,y) is the coordinate of the projectile at any time t seconds. Both x&y are measured in meters.

a) find the range of the projectile

b) will the projectile hit the too of the cliff 50m high?

c) what is the coordinate of the projectile at t=6seconds


1
Expert's answer
2021-02-15T05:43:13-0500

Solution:


x = v0x t

y = v0y t - "\frac{gt^2}{2}"

eliminating t on get the trajectory:


"y=tan(\alpha)x-[\frac{g}{2v_0x^2}]x^2"


then

"tan(\alpha)= \sqrt{\smash[b]{\frac{3}{3}}} \rightarrow\; \alpha=30 ^o"


"\dfrac{g}{2v_ox^2}=\dfrac{4}{3}\rightarrow v_ox=1.92(\frac{m}{s})\rightarrow v_o=2.21(\frac{m}{s})"


so:

a)

"x=\dfrac{v_o^2sin(2 \alpha)}{g}=0.43m"

b)

"h=\dfrac{v_oy^2}{2g}=0.06m"


c)

fly time tf is

tf = "\dfrac{2v_oy}{g}=0.23s"

then being t = 6 s > tf

x = vox tf = 0.43 m

y = 0